# Weight-Space Linear Recurrent Neural Networks

### v1

@@ B.1 Datasets
Problem Dataset # Samples $N$ Seq. Length $T$ Context Length $L$ Features $D_{x}$
2D Images MNIST 60,000 784 (varies) 1
@@
Both MNIST LeCun et al. [1998] and Fashion MNIST Xiao et al. [2017] datasets were loaded using the well-known PyTorch interface Paszke et al. [2017]. The values were then normalized so that pixel values ranged $[-1,1]$. The CelebA dataset Liu et al. [2015] was loaded using the API from Nzoyem et al. [2024] itself inspired by Zintgraf et al. [2019]. Training was performed on the train sets (see attributes in Table 4), with validation and testing on the predefined test sets.
@@
For the electricity data Zhou et al. [2021], we further normalized the preprocessed data from TSLib to place all values in the range $[-1,1]$ in order to facilitate learning dynamics. We did not use the predefined “test” set because of its 144 steps-long forecast window, which is much longer than the 96 steps all models saw during training [^7] Forecasting experiments with inference beyond the training regime, i.e., $\tau>1$, are reserved for future work.. Consequently, we used the “validation” set to evaluate our models as well as all the baselines.
@@
On the UEA dataset Bagnall et al. [2018], we follow the procedure from Walker et al. [2024] and reuse the same dataset. (We note that this exact experimental protocol was recently observed in Rusch and Rus [2024]). We extracted the necessary scripts for reproducibility and provide them as part of our code under appropriate license.
@@
Continuous autonomous dynamical systems can be conceptualized as multivariate time series governed by a deterministic vector field $(\mathbf{x}_{\tau},p)\mapsto\dot{\mathbf{x}}_{\tau}$, with $p$ encompassing physical parameters affecting the dynamics. Given an initial condition $\mathbf{x}_{0}$, one can systematically simulate and subsample the trajectory $\mathbf{x}_{0:T}$. To complement our description in Section 3.2, we provide the vector field used for each dataset in Table 5. We summarize their physical parameter ranges in Table 6. All trajectories are obtained with SciPy’s ‘RK45’ adaptive time-stepping numerical integrator Virtanen et al. [2020]. The five training data splits of the SINE are “Tiny”, “Small”, “Medium”, “Large”, “Huge”, with respectively 1, 10, 100, 1k, and 10k samples. All datasets are normalized and placed within the $[-1,1]$ which is calculated using train set statistics. Comprehensive data generation scripts with physical parameters for all four datasets are provided in our code.
@@
The Spirals dataset is an additional dynamical system dataset for binary classification tasks. The training data consists of 10,000 samples, where each sample is a spiral trajectory represented as a sequence of 64 2D points ($x,y$ coordinates). Half of the dataset contains clockwise spirals (labeled as 0), while the other half contains counterclockwise spirals (labeled as 1). The spirals are generated using sine and cosine functions with random phase offsets, and the amplitude decreases with time to create the spiral effect. This dataset serves as a powerful test case for dynamics classification; it was inspired by Kidger and Garcia [2021] [^8] An example usage can be found at https://docs.kidger.site/diffrax/examples/neural_cde/, with a few samples visualized in Figure 9.
Figure 9: Visualization of a few samples of the Spirals datasets. For each sample, we plot the $y$ and the $x$ sequences of coordinates against each other to observe the (counter-)clockwise direction.
B.2 Baselines
@@
We consider two powerful RNN baselines including the Gated Recurrent Unit (GRU) Cho et al. [2014] and the Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber [1997]. Both are trained in recurrent AR mode for forecasting problems and recurrent non-AR mode for classification. Depending on the experiment, we vary their hidden size to match the total parameter count of WARP. The remainder of the experiment details such as training procedure are presented in Appendix C. We attach both implementations, using Equinox Kidger and Garcia [2021], to our code.
@@
We consider the Time Series Transformer from HuggingFace Jain [2022]. This baseline provides a SoTA baseline, leveraging one of the most transformative sequence mixing processes to date: Attention. The specific model used was the TimeSeriesTransformerForPrediction, which adds a distribution head on top of the vanilla encoder-decoder Transformer Vaswani et al. [2017]. This means that the model learns a distribution, from which we take the mean to be used for time-series forecasting. The next token prediction is obtained by randomly sampling a window of context length $L$ plus prediction length $T-L$ from the target time series. This prediction window is subsequently masked for the next token prediction task.
@@
The ConvCNP Gordon et al. [2020] is an encoder-based meta-learning approach which doesn’t require gradients in order to adapt to novel scenarios. The ConvCNP is trained for on-the-grid image completion with 100 random shots (time steps. We adapt its data loading process to allow the ConvCNP to operate on raster-scan-ordered pixels at test time.
@@
We use the powerful Structured State Space Model S4 with the implementation of Rush and Karamcheti [2022]. We particularly apply it to the MNIST experiment, where the goal is to forecast a contiguous range of future predictions given a range of past contexts. To that end, we simply concatenate the entire context with a sequence of masks set to the length of the forecast window. This input is a single sequence of length $T$ that is run through the deep S4 model, which maps to an output of length $T$. We then use the last $T-L$ tokens as the forecasted predictions.
@@
NCDEs or (Neural CDEs) Kidger et al. [2020] provide a continuous-time framework for processing irregularly-sampled time series by interpreting the data as a continuous path. By using the path as a control for a neural differential equation, NCDEs can naturally handle missing data and irregular sampling. The continuous nature of NCDEs makes them a strong baseline exclusively for classification tasks.
@@
Mamba, S6, Log-CDE, NRDE, NCDE, LRU were all reported from Walker et al. [2024], where we direct the reader for further details. We reused the results and the conclusion from that work, as was done by Rusch and Rus [2024].
B.3 Metrics
@@
The (BPD) is used to evaluate the quality of generative models, particularly for images. It quantifies how many bits are needed on average to encode each dimension (e.g., pixel) of the data, with lower BPD values indicating a better model. The BPD is derived from the negative log-likelihood (NLL) of the data under the model’s predicted distribution. For a given ground truth pixel value $\mathbf{y}_{t}$ and its corresponding predicted mean $\hat{\mathbf{y}}_{t}$ and standard deviation $\hat{\boldsymbol{\sigma}_{t}}$, the overall NLL over the image is calculated as:
$\text{NLL}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{2}\log(2\pi\hat{\boldsymbol{\sigma}_{t}}^{2})+\frac{1}{2}\frac{(\mathbf{y}_{t}-\hat{\mathbf{y}}_{t})^{2}}{\hat{\boldsymbol{\sigma}_{t}}^{2}}.$
@@
Appendix C Experimental Details
@@
Although specific details may vary depending on the problem, the root network is consistently chosen as a MLP for all problem sets in this paper. Given the quadratic memory cost $O(D_{\theta}^{2})$, we can vary its layers to balance batch size with capacity. Image completion and forecasting problems use the ReLU activation Agarap [2018], while smooth dynamical system reconstruction use the Swish Ramachandran et al. [2017]. Complete details on the root network are given in Table 7. The initial hypernetwork $\phi$ —used for all problems except Image Completion, MSD, and LV— is made up of two hidden layers of width $\nicefrac{{h_{\text{in}}}}{{3}}+\nicefrac{{2h_{\text{out}}}}{{3}}$ and $\nicefrac{{2h_{\text{in}}}}{{3}}+\nicefrac{{h_{\text{out}}}}{{3}}$ neurons, respectively. The positive integers $h_{\text{in}}$ and $h_{\text{out}}=D_{\theta}$ are the number of input and output neurons respectively.
WARP-Phys setup.
For the SINE experiment, the root network predicts the phase $\hat{\varphi}$ to feed into the sinusoid $\tau\mapsto\sin(2\pi\tau+\hat{\varphi})$. For the challenging MSD and MSD-Zero problems, we embed knowledge of the general analytical solution and the initial condition with $\tau\mapsto E(\tau)\mathbf{x}_{0}$, where $E(\cdot)\in\mathbb{R}^{2\times 2}$ with its four coefficients predicted by the root network, and $\mathbf{x}_{0}$ is known throughout. $E(\tau)$ is viewed as the exponential of $\tau A$, where $A$ is the constant matrix characterizing the mass-spring-damper dynamics: $A=\big{(}\begin{smallmatrix}0&1\\ -k/m&-c/m\end{smallmatrix}\big{)}$ Nzoyem and Labbe [2021]; its poor conditioning — a consequence of the large parameter scales and ranges detailed in Table 6 — would destabilize the training if $A$ was directly learned. We note that stronger levels of physics may be embedded into the root network: predicting rescaled time-invariant constants $(m,k,c)$, parameterizing the signal as damped sinusoids, eigen-decomposition, etc. The physics-informed strategy we present is the one that produced the biggest improvement over WARP in our experiments.
Optimization & Core baselines.
Our WARP framework (along with our custom GRU, LSTM, ConvCNP, and Neural CDE) is implemented with the JAX framework Bradbury et al. [2018] and its ecosystem: Equinox for neural network definitions Kidger and Garcia [2021], and Optax for optimization DeepMind et al. [2020]. We use the AdaBelief optimizer Zhuang et al. [2020], and we clip the gradient norm with $g_{\text{lim}}=10^{-7}$. We apply the “reduce on plateau” rule where the learning rate is divided by 2 if the average loss [^9] The average being calculated over 50 iterations. doesn’t evolve after 20 epochs. For most problems, we set the initial learning rate at $10^{-5}$. All GRU and LSTM models have a single layer to match WARP. We tweak their hidden size rather than the number of layers in order to increase or reduce parameter count, thus keeping in check the complexity of the models under consideration.
TST and S4 baselines setup.
As for the TST model for forecasting, it is implemented in PyTorch Paszke et al. [2017] and uses the AdamW optimizer with a constant learning rate of $6\times 10^{-4}$, beta values of $0.9$ and $0.95$ and weight decay coefficient of $10^{-1}$. For the S4 model, we used 6 layers, each with a hidden state of size 512, a batch size of 50, a learning rate of $10^{-3}$, and the traditional weight decay with coefficient $0.05$. Like Rush and Karamcheti [2022], we used prenormalization, but we did not use dropout. These hyperparameters were selected based on the validation set when available, and the test set if not (e.g., all synthetic toy problems).
Hardware.
Table 7: Root MLP configurations for each dataset for each problem.
@@
2D Images MNIST 24 3 ReLU
@@
Dynamical Systems MSD 48 3 Swish
@@
C.1 Image Completion
On these problems, since trained with the NLL loss from Equation 3, we use a dynamic tanh activation function on the mean prediction, with $(a,b,\alpha,\beta)$ initialized as $(1,0,1,0)$. Only experiments run on MNIST and Fashion MNIST use weight clipping, while CelebA does not [^10] Apart from (Fashion) MNIST, no other problem in this work used weight clipping.. For CelebA, we use $\sigma_{\text{lim}}=10^{-4}$ (see Equation 2) while we find the unusually large $\sigma_{\text{lim}}=0.5$ suitable for MNIST and Fashion MNIST. We compare WARP to the MNIST baselines roughly at the same parameter counts: GRU (1.694 M), LSTM (1.696 M), S4 (1.717 M), and WARP (1.687 M). We train for 200 and 250 epochs in batches of 640 and 1256 for (Fashion) MNIST and CelebA respectively. We apply the recurrent AR mode with $p_{\text{forcing}}=0.15$, while directly feeding the mean prediction back into the recurrence (i.e., the reparameterization trick is disabled both during training and inference).
C.2 Image classification
@@
C.3 Energy forecasting
For these experiments, all models share identical architectures across the four datasets (ETT-h1, ETT-h2, ETT-m1, ETT-m2). The learning rate differs between hourly and minute-level datasets: $10^{-5}$ for h1/h2 and $10^{-4}$ for m1/m2. For hourly datasets, we train for 500 epochs, while minute-level datasets require only 250 epochs (corresponding to roughly 1.5 hours of training). All models are trained with batch size 3600 in autoregressive mode with stochastic sampling (non-AR), and $p_{\text{forcing}}=0.25$. No final activation is applied to the root network’s mean output, while the typical positivity-enforcing from Equation 2 is applied to the standard deviation with $\sigma_{\text{lim}}=10^{-4}$.
C.4 Dynamical System Reconstruction
For the dynamical system reconstruction tasks, since uncertainties are not required, models are trained without NLL loss (i.e., with the MSE loss defined in Equation 3). Consistent across all dynamical system experiments, no weight clipping is employed, and the predictions are enforced in the range $[-1,1]$ with a unit-initialized dynamic tanh. All losses are computed on the normalized test set.

### v2

@@ C.1 Datasets
Problem Dataset # Samples NN Seq. Length $T$ Context Length $L$ # Features $D_{x}$
2D Images
MNIST 60,000 784 (varies) 1
@@
Both MNIST [58] and Fashion MNIST [96] datasets were loaded using the well-known PyTorch interface [76]. The values were then normalised so that pixel values ranged [1,1][-1,1]. The CelebA dataset [61] was loaded using the API from [72] itself inspired by [108]. Training was performed on the train sets (see attributes in Table 5), with validation and testing on the predefined test sets.
@@
For the electricity data [105], we further normalised the preprocessed data from TSLib to place all values in the range [1,1][-1,1] in order to facilitate learning dynamics. We did not use the predefined “test” set because of its 144-step-long forecast window, which is much longer than the 96 steps all models saw during training. Consequently, we used the “validation” set to evaluate our models as well as all the baselines.
@@
On the UEA dataset [8], we follow the procedure from [93] and reuse the same dataset. (We note that this exact experimental protocol was recently observed in [78]). We extracted the necessary scripts for reproducibility and provide them as part of our code under appropriate license.
@@
Continuous autonomous dynamical systems can be conceptualised as multivariate time series governed by a deterministic vector field $(\mathbf{x}_{\tau},p)\mapsto\dot{\mathbf{x}}_{\tau}$, with $p$ encompassing physical parameters affecting the dynamics. Given an initial condition $\mathbf{x}_{0}$, one can systematically simulate and subsample the trajectory $\mathbf{x}_{0:T}$. To complement our description in Section 3.2, we provide the vector field used for each dataset in Table 6. We summarise their physical parameter ranges in Table 7. All trajectories are obtained with SciPy’s ‘RK45’ adaptive time-stepping numerical integrator [90]. The five training data splits of the SINE are “Tiny”, “Small”, “Medium”, “Large”, “Huge”, with respectively 1, 10, 100, 1k, and 10k samples. All datasets are normalised and placed within [1,1][-1,1], which is calculated using the train set statistics. Comprehensive data generation scripts with physical parameters for all four datasets are provided in our code.
@@
The Spirals dataset is an additional dynamical system dataset for binary classification tasks. The training data consists of 10,000 samples, where each sample is a spiral trajectory represented as a sequence of 64 2D points ($x,y$ coordinates). Half of the dataset contains clockwise spirals (labelled as 0), while the other half contains counterclockwise spirals (labelled as 1). The spirals are generated using sine and cosine functions with random phase offsets, and the amplitude decreases with time to create the spiral effect. This dataset serves as a powerful test case for dynamics classification; it was inspired by [52] [^8] An example usage can be found at https://docs.kidger.site/diffrax/examples/neural_cde/, with a few samples visualised in Fig. 7.
Figure 7: Visualisation of a few samples of the Spirals datasets. For each sample, we plot the $y$ and the $x$ sequences of coordinates against each other to observe the (counter-)clockwise direction.
C.2 Baselines
@@
We consider two powerful RNN baselines including the Gated Recurrent Unit (GRU) [18] and the Long Short-Term Memory (LSTM) [43]. Both are trained in recurrent AR mode for forecasting problems and recurrent non-AR mode for classification. Both are unidirectional and have a single layer to match our WARP model. Depending on the experiment, we vary their hidden size to match the total parameter count of WARP. The remainder of the experiment details such as training procedure are presented in Appendix D. We attach both implementations, using Equinox [52], to our code.
@@
We consider the Time Series Transformer from HuggingFace [45]. This baseline provides a SoTA baseline, leveraging one of the most transformative sequence mixing processes to date: Attention. The specific model used was the TimeSeriesTransformerForPrediction, which adds a distribution head on top of the vanilla encoder-decoder Transformer [89]. This means that the model learns a distribution from which we take the mean to be used for time-series forecasting. The next token prediction is obtained by randomly sampling a window of context length $L$ plus prediction length TLT-L from the target time series. This prediction window is subsequently masked for the next token prediction task.
@@
The ConvCNP [30] is an encoder-based meta-learning approach that doesn’t require gradients in order to adapt to novel scenarios. The ConvCNP is trained for on-the-grid image completion with 100 random shots (time steps. We adapt the data loading process to allow the ConvCNP to operate on raster-scan-ordered pixels at test time.
@@
We use the powerful Structured State Space Model S4 with the implementation of [79]. We particularly apply it to the MNIST experiment, where the goal is to forecast a contiguous range of future predictions given a range of past contexts. To that end, we simply concatenate the entire context with a sequence of masks set to the length of the forecast window. This input is a single sequence of length $T$ that is run through the deep S4 model, which maps to an output of length $T$. We then use the last TLT-L tokens as the forecasted predictions.
@@
NCDEs or (Neural CDEs) [53] provide a continuous-time framework for processing irregularly-sampled time series by interpreting the data as a continuous path. By using the path as a control for a neural differential equation, NCDEs can naturally handle missing data and irregular sampling. The continuous nature of NCDEs makes them a strong baseline exclusively for classification tasks.
@@
Mamba, S6, Log-CDE, NRDE, NCDE, LRU were all reported from [93], where we direct the reader for further details. We reused the results and the conclusion from that work, as was done by Rusch & Rus [78].
C.3 Metrics
@@
The (BPD) is used to evaluate the quality of generative models, particularly for images. It quantifies how many bits are needed on average to encode each dimension (e.g., pixel) of the data, with lower BPD values indicating a better model. The BPD is derived from the negative log-likelihood (NLL) of the data under the model’s predicted distribution. For a given ground truth pixel value $\mathbf{y}_{t}$ and its corresponding predicted mean $\hat{\mathbf{y}}_{t}$ and standard deviation $\hat{\bm{\sigma}_{t}}$, the overall NLL over the image is calculated as:
$\text{NLL}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{2}\log(2\pi\hat{\bm{\sigma}_{t}}^{2})+\frac{1}{2}\frac{(\mathbf{y}_{t}-\hat{\mathbf{y}}_{t})^{2}}{\hat{\bm{\sigma}_{t}}^{2}}.$
@@
Appendix D Experimental Details
@@
Although specific details may vary depending on the problem, the root network is consistently chosen as an MLP for all problem sets in this paper. Given the quadratic memory cost $O(D_{\theta}^{2})$, we can vary its layers to balance batch size with capacity. Image completion and forecasting problems use the ReLU activation [1], while smooth dynamical system reconstruction uses the Swish [77]. Complete details on the root network are given in Table 8. The initial hypernetwork $\phi$ —used for all problems except Image Completion, MSD, and LV— is made up of two hidden layers of width $\nicefrac{{h_{\text{in}}}}{{3}}+\nicefrac{{2h_{\text{out}}}}{{3}}$ and $\nicefrac{{2h_{\text{in}}}}{{3}}+\nicefrac{{h_{\text{out}}}}{{3}}$ neurons, respectively. The positive integers $h_{\text{in}}$ and $h_{\text{out}}=D_{\theta}$ are the number of input and output neurons, respectively.
Table 8: Root MLP configurations for the datasets in each problem.
@@
2D Images
MNIST 24 3 ReLU
@@
Dynamical
Systems
MSD 48 3 Swish
@@
WARP-Phys setup.
[…]
Optimisation & Core baselines.
Our WARP framework (along with our custom GRU, LSTM, ConvCNP, and Neural CDE) is implemented with the JAX framework [14] and its ecosystem: Equinox for neural network definitions [52], and Optax for optimisation [24]. We use the AdaBelief optimiser [107], and we clip the gradient norm with $g_{\text{lim}}=10^{-7}$. We apply the “reduce on plateau” rule where the learning rate is divided by 2 if the average loss [^9] The average being calculated over 50 iterations. doesn’t evolve after 20 epochs. For most problems, we set the initial learning rate at $10^{-5}$. All GRU and LSTM models have a single layer to match WARP. We tweak their hidden size rather than the number of layers in order to increase or reduce parameter count, thus keeping in check the complexity of the models under consideration (see Table 9).
@@
TST and S4 baselines setup.
As for the TST model for forecasting, it is implemented in PyTorch [76] and uses the AdamW optimiser with a constant learning rate of $6\times 10^{-4}$, beta values of $0.9$ and $0.95$ and weight decay coefficient of $10^{-1}$. For the S4 model, we used 6 layers, each with a hidden state of size 512, a batch size of 50, a learning rate of $10^{-3}$, and the traditional weight decay with coefficient $0.05$. Like [79], we used prenormalisation, but we did not use dropout. These hyperparameters were selected based on the validation set when available, and the test set if not (e.g., all synthetic toy problems).
Hardware.
D.1 Image Completion
On these problems, since trained with the NLL loss from Eq. 2, we use a dynamic tanh activation function on the mean prediction, with $(a,b,\alpha,\beta)$ initialised as $(1,0,1,0)$. Only experiments run on MNIST and Fashion MNIST use weight clipping, while CelebA does not [^10] Apart from (Fashion) MNIST, no other problem in this work used weight clipping.. For CelebA, we use $\sigma_{\text{lim}}=10^{-4}$ (see Section 2.2) while we find the unusually large $\sigma_{\text{lim}}=0.5$ suitable for MNIST and Fashion MNIST. We compare WARP to the MNIST baselines roughly at the same parameter counts: GRU (1.694 M), LSTM (1.696 M), S4 (1.717 M), and WARP (1.687 M). We train for 200 and 250 epochs in batches of 640 and 1256 for (Fashion) MNIST and CelebA, respectively. We apply the recurrent AR mode with $p_{\text{forcing}}=0.15$, while directly feeding the mean prediction back into the recurrence (i.e., the reparametrisation trick is disabled both during training and inference).
As inputs to the root network, while the normalised pixel coordinates are better suited for this task, we report our state-of-the-art results using the normalised time $\tau=1/(T-1)$. In fact, all results presented in the paper only use the normalised time coordinate system, except for the time series classification on the UEA dataset presented below.
D.2 Image Classification
@@
D.3 Energy Forecasting
For these experiments, all models share identical architectures across the four datasets (ETT-h1, ETT-h2, ETT-m1, ETT-m2). The learning rate differs between hourly and minute-level datasets: $10^{-5}$ for h1/h2 and $10^{-4}$ for m1/m2. For hourly datasets, we train for 500 epochs, while minute-level datasets require only 250 epochs (corresponding to roughly 1.5 hours of training). All models are trained with batch size 3600 in autoregressive mode with stochastic sampling (non-AR), and $p_{\text{forcing}}=0.25$. No final activation is applied to the root network’s mean output, while the typical positivity-enforcing from Section 2.2 is applied to the standard deviation with $\sigma_{\text{lim}}=10^{-4}$.
D.4 Traffic Flow Forecasting
Our model architecture disregards the explicit spatial connectivity provided in the PEMS08 dataset [85]. Instead, we consider the features from all nodes independently, creating a flattened feature vector for each time step. This results in an input of shape (12, 510), where 12 is the number of historical time steps and 510 represents the 170 nodes, each with 3 features. Before the input sequence is used in the linear recurrence, its features are transformed by a 1D-convolution with 510 input channels, 4080 output channels, and a kernel length of 36. The model is trained to predict a single feature per node for the future 12 time steps.
D.5 In-Context Learning
The setting is the elegant in-context learning setting developed by [99], where the goal is learn the linear mapping between several key-value pairs. The keys $\{\mathbf{x}_{i}\}_{i=1,\ldots,N}$ are vectors of dimension $D_{x}-1$, and the values $\{y_{i}\}_{i=1,\ldots,N}$ are scalar, both concatenated to form a state of dimension $D_{x}$. A final query key is given, and the model must predict its corresponding value (substituted by 0 in the input sequence).
Importantly, to retain consistency across the literature, we preserve the notations from [99], even though they conflict with those established in our problem setting in Section 2.1. To revert back to our original setting, one can replace the existing inputs with $\mathbf{x}_{t}\triangleq\text{concat}(\mathbf{x}_{t+1},y_{t+1})$, for $t=0,\ldots,T-2$; and $\mathbf{x}_{T-1}\triangleq\text{concat}(\mathbf{x}_{q},0)$. As for the outputs, $\mathbf{y}_{t}\triangleq y_{t+1}$, for $t=0,\ldots,T-2$; and $\mathbf{y}_{T-1}\triangleq y_{q}$.
D.6 Dynamical System Reconstruction
For the dynamical system reconstruction tasks, since uncertainties are not required, models are trained without NLL loss (i.e., with the MSE loss defined in Eq. 2). Consistent across all dynamical system experiments, no weight clipping is employed, and the predictions are enforced in the range [1,1][-1,1] with a unit-initialised dynamic tanh. All losses are computed on the normalised test set.